A 14 meter ladder is leaning against a building. The bottom of the ladder is 7 meters from the building. How high is the top of the ladder?
Answer by checkley77(12844)

7^2+x^2=14^2
49+x^2=196
x^2=196-49
x^2=147
x=sqrt147
x=12.124 meters is the height of the ladder against the wall.